Solve. (x – 5)^2 = 7

Answer by stanbon(75887)

Solve. (x – 5)^2 = 7
Take the square root of both sides to get:
x-5 = sqrt7 or x-5 = -sqrt7
x=5+sqrt7 or x=5-sqrt7
=================
Cheers,
Stan H.